two numbers differ by 12
a. if the smaller number is s, then the larger number is?
b. if the larger number is l, then the smaller number is?

Answer by mananth(16949)


let the numbers be x & y
x-y=12
y=s
x-s=12
add s to both sides
x-s+s=12+s
x=12+s the larger number
....
...
larger number is 1
x=1
1-y=12
add y to both sides
1-y+y=12+y
1=12+y
add -12 to both sides
1-12=12-12+y
y=-11 the smaller number is -11